OpenStudy (rock_mit182):

A freight train leaves the train station 44 hourshours before a passenger train. The two trains are traveling in the same direction on parallel tracks. If the rate of the passenger train is 40 mph faster than the freight​ train, how fast is each train traveling if the passenger train passes the freight train in 66 ​hours? How fast is the freight train​ traveling? _____ mph

OpenStudy (rock_mit182):

A freight train leaves the train station 4 hours *

OpenStudy (rock_mit182):

how fast is each train traveling if the passenger train passes the freight train in 6 ​hours?

OpenStudy (templeguy):

I'll answer yours if you answer mine

OpenStudy (templeguy):

show you what?

OpenStudy (templeguy):

Can you answer this? What was the joke from 17:51 - 18:03 in this vid and why did she spray the air freshener https://www.youtube.com/watch?v=Rhn5uobZAyQ?

OpenStudy (rock_mit182):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (rock_mit182):

hey bro walk me through please

OpenStudy (3mar):

3 minutes, please!

OpenStudy (sshayer):

let the speed of freight train =v m/p speed of passenger train=v+40 m/h then V*(44+66)=(v+40)*66 find v and then v+40

OpenStudy (rock_mit182):

mm ok i can do that

OpenStudy (rock_mit182):

with the correct data would be : V (4 +6) = (v +40) 6

OpenStudy (rock_mit182):

v =60 60 + 40 =100 How fast is the freight train​ traveling ?

OpenStudy (rock_mit182):

thanks for your help

OpenStudy (sshayer):

110v=66v+66*40 110v-66v=66*40 44v=66*40 \[v=\frac{ 66*40 }{ 44 }=?\]